The aquarium charges $16.00 for admission and $2.50 for each ride ticket. Mountain rides charges $25.00 for admission and $0.25 for each ride ticket. For what number of ride tickets will the cost be the same at either venue? What is the cost?
Found 2 solutions by solver91311, greenestamps:
Answer by solver91311(24713)





Solve for , Calculate


John

My calculator said it, I believe it, that settles it


Answer by greenestamps(13200)



Without algebra....

The difference in the admission cost is $9.00; the difference in the cost of each ride is $2.25.

The number of rides needed to match the difference in admission costs is 9.00/2.25 = 4.

The cost of each venue for 4 rides is 16+4(2.50) = 16+10 = 26 or 25+4(0.25) = 25+1 = 26.